### v2

@@ 5 Transformer model
  • Let $Z\in\mathbb{R}^{d\times\omega}$. This is the embedding matrix. Its $\gamma$th column corresponds to the embedding of token $\gamma$ in $\mathbb{R}^{d}$. We call $d$ the embedding dimension.
@@
  • Let $W\in\mathbb{R}^{d\times m},b\in\mathbb{R}^{m},V\in\mathbb{R}^{m\times\omega}$. These are the hidden matrix, bias vector, and output matrix respectively. We call $m$ the number of neurons.
@@
$\begin{split}&h_{1,\theta}:\cup_{t=1}^{\infty}[\omega]^{t}\to\cup_{t=1}^{\infty}\mathbb{R}^{d\times t}:\alpha\mapsto Z[:,\alpha]+U[:,1:|\alpha|],\\ &h_{2,\theta}:\cup_{t=1}^{\infty}\mathbb{R}^{d\times t}\to\mathbb{R}^{d}:X\mapsto W_{0}^{\top}\text{vec}\left(\left[W_{3,r}^{\top}X\phi\left(X^{\top}W_{1,r}W_{2,r}^{\top}X[:,-1]\right)\right]_{r\in[m_{0}]}\right),\\ &h_{3,\theta}:\mathbb{R}^{d}\to\mathbb{R}^{\omega}:x\mapsto V^{\top}\phi\left(W^{\top}x+b\right),\\ \text{and }&h_{\theta}\coloneqq\phi\circ h_{3,\theta}\circ h_{2,\theta}\circ h_{1,\theta}\end{split}$ (6)
where $-1$ denotes the last index and $\phi$ is the softmax function. We think of the value $h_{\theta}(\leavevmode\nobreak\ )$ as an additional $\omega-1$ model parameters. The first three functions are the embedding, self-attention, and FNN sub-layers respectively of the one-layer transformer $h_{\theta}$. Note that we do not include the skip-connection $+X[:,-1]$ in the self-attention sub-layer. This was omitted for simplicity, but its inclusion does not affect Theorems 6.1 and 6.5, which follow by almost the same proof.
@@ 6 Transformer next-token prediction capacity
In order to lower bound the next-token prediction capacity of the transformer model, we first reduce to the special case of $d=d_{1}=d_{0}=m_{0}=1$ and $W_{1,1}=W_{2,1}=W_{3,1}=W_{0}=1$. Thus, the model becomes the following composition of mappings:
@@
where $f(z,u,\alpha)=\langle x,\phi(xx_{|\alpha|})\rangle$ with $x=z[\alpha]+u[1:|\alpha|]$. Assume $\psi:\mathbb{R}\to\mathbb{R}$ is real analytic at some $\eta\in\mathbb{R}$ and not a polynomial there. Assume $m\geq n$. We will show that this simplified model can interpolate arbitrary data sets of size $n$. Then we will use this to show that the full model can interpolate arbitrary data sets of size $n$.
@@
To show the former, let $\alpha^{1},\ldots,\alpha^{n}\in\cup_{t=1}^{\infty}[\omega]^{t}$ and $Y^{\prime}\in(\mathrm{ri}\Delta^{\omega-1})^{n}$. Then there exists $Y\in\mathbb{R}^{\omega\times n}$ such that $\phi(Y)=Y^{\prime}$ with $\phi$ applied column-wise. We will construct an interpolating model. First, set $b=\eta\mathbbm{1}_{m}$. Second, define $T=\max_{i\in[n]}|\alpha^{i}|$ and sample $(z,u)\in\mathbb{R}^{\omega}\times\mathbb{R}^{T}$ from any continuous probability distribution. Let $\hat{x}_{1},\ldots,\hat{x}_{n}$ be the corresponding outputs of the first mapping. Then the matrix of outputs of the composition of mappings is $\phi(V^{\top}\psi(w[\hat{x}_{i}]_{i\in[n]}^{\top}+b\mathbbm{1}_{n}^{\top}))$. In Theorem 6.1, we prove that the $\hat{x}_{i}$ are nonzero and distinct. In Theorem 6.4, we prove that this implies $\psi(w[\hat{x}_{i}]_{i\in[n]}^{\top}+b\mathbbm{1}_{n}^{\top})$ has full rank for generic $w$ in an open set around zero. So, as our third step in the construction, sample $w\in\mathbb{R}^{m}$ from any continuous probability distribution that has support in the open set. Then, $\psi(w[\hat{x}_{i}]_{i\in[n]}^{\top}+b\mathbbm{1}_{n}^{\top})$ has rank $n$ so, as our fourth and final step in the construction, solve for $V$ in $Y=V^{\top}\psi(w[\hat{x}_{i}]_{i\in[n]}^{\top}+b\mathbbm{1}_{n}^{\top})$ as a system of linear equations.
@@
Thus, the simplified model is able to interpolate arbitrary data sets as long as $m\geq n$. To see how to reduce the full model with $d,m_{0},d_{0}\in\mathbb{N}$ and $d_{r}\in\mathbb{N}\leavevmode\nobreak\ \forall r\in[m_{0}]$ to the simplified model, set $Z=\mathbbm{1}_{d}z^{\top}/\sqrt{d},U=\mathbbm{1}_{d}u^{\top}/\sqrt{d},W_{1,r}=W_{2,r}=\mathbbm{1}_{d}\mathbbm{1}_{d_{r}}^{\top}/\sqrt{dd_{r}}\leavevmode\nobreak\ \forall r\in[m_{0}],W_{3,r}=\mathbbm{1}_{d}\mathbbm{1}_{d_{0}}^{\top}/\sqrt{dd_{0}}\leavevmode\nobreak\ \forall r\in[m_{0}],W_{0}=\mathbbm{1}_{m_{0}d_{0}}\mathbbm{1}_{d}^{\top}/(m_{0}\sqrt{dd_{0}}),W=\mathbbm{1}_{d}w^{\top}/\sqrt{d}$. In other words, token $\gamma\in[\omega]$ is embedded as $z_{\gamma}\mathbbm{1}_{d}/\sqrt{d}$ and the $t$th position vector is $u_{t}\mathbbm{1}_{d}/\sqrt{d}$. So, the inputs to the self-attention sub-layer are sequences of vectors coming from a one-dimensional subspace, and so it makes sense for the self-attention matrices to be constant matrices. Finally, the $k$th hidden neuron of the FNN sub-layer has weight vector $w_{k}\mathbbm{1}_{d}/\sqrt{d}$. By reducing to the simplified model in this way, and applying Theorems 6.1 and 6.4, we are able to lower bound the next-token prediction capacity of the one-layer multi-head decoder-only transformer model by $m$ in Theorem 6.5.
@@ 6.1 Injectivity of self-attention
Theorem 6.1.
Let $\omega,T\in\mathbb{N}$. Let $\phi$ be the softmax function. Define
$f:\leavevmode\nobreak$ $\mathbb{R}^{\omega}\times\mathbb{R}^{T}\times\cup_{t=1}^{T}[\omega]^{t}\to\mathbb{R}$
@@
First, let $A$ denote $\cup_{t=1}^{T}[\omega]^{t}$. Then, note that, for all $z\in\mathbb{R}^{\omega},u\in\mathbb{R}^{T},\tau\in[T],\alpha\in[\omega]^{\tau}$,
@@
is real analytic. Note that the set of $(z,u)$ for which $f(z,u,\cdot)$ does not map injectively into \$\0\mathbb{R}\backslash\{0\}$ is precisely the zero set of $g$. Thus, the result follows if and only if $g$ is nontrivial if and only if (1) each $f(\cdot,\cdot,\alpha)$ is nontrivial and (2) each $f(\cdot,\cdot,\alpha)-f(\cdot,\cdot,\beta)$ is nontrivial. To prove (1), observe that $f(\mathbbm{1}_{\omega},0,\alpha)=1\leavevmode\nobreak\ \forall\alpha\in A$. Now, we will go about proving (2). First, let $\alpha,\beta\in A$. We want to show that $\alpha\neq\beta$ implies $f(\cdot,\cdot,\alpha)-f(\cdot,\cdot,\beta)$ is nontrivial. We will prove the contrapositive. Suppose $f(\cdot,\cdot,\alpha)-f(\cdot,\cdot,\beta)\equiv 0$. First, we will show that $\alpha$ and $\beta$ have the same length, then we will show that each entry of $\alpha-\beta$ is zero.
@@
Let $\tau=|\alpha|=|\beta|$. Let $s\in[\tau]$. To show that $\alpha_{s}=\beta_{s}$, set $z=(1,\ldots,\omega)$ and $u=\omega e_{s}$, and define the polynomials
$p_{\gamma}=\sum_{t=1}^{\tau}(\gamma_{t}+\omega\delta_{t,s})x^{(\gamma_{t}+\omega\delta_{t,s})(\gamma_{\tau}+\omega\delta_{\tau,s})}\text{ and}q_{\gamma}=\sum_{t=1}^{\tau}x^{(\gamma_{t}+\omega\delta_{t,s})(\gamma_{\tau}+\omega\delta_{\tau,s})}\leavevmode\nobreak\ \forall\gamma\in\{\alpha,\beta\}.$
Note that the monomial $x^{(\alpha_{s}+\omega)(\alpha_{\tau}+\omega\delta_{\tau,s})+(\beta_{s}+\omega)(\beta_{\tau}+\omega\delta_{\tau,s})}$ has coefficient $\alpha_{s}-\beta_{s}$ in $p_{\alpha}q_{\beta}-p_{\beta}q_{\alpha}$. Also, $f(z,u,\gamma)=p_{\gamma}(\exp(1))/q_{\gamma}(\exp(1))\leavevmode\nobreak\ \forall\gamma\in\{\alpha,\beta\}$. So, again, $p_{\alpha}q_{\beta}-p_{\beta}q_{\alpha}\equiv 0$. Thus, $\alpha_{s}=\beta_{s}$. Since this was for arbitrary $s\in[\tau]$, we get that $\alpha=\beta$, proving the theorem. ∎
@@
$\frac{\sum_{t=1}^{\tau}(z_{\alpha_{t}}+u_{t})\exp((z_{\alpha_{t}}+u_{t})(z_{\alpha_{\tau}}+u_{\tau}))}{\sum_{t=1}^{\tau}\exp((z_{\alpha_{t}}+u_{t})(z_{\alpha_{\tau}}+u_{\tau}))}=\frac{\sum_{t=1}^{\tau}(z_{\beta_{t}}+u_{t})\exp((z_{\beta_{t}}+u_{t})(z_{\beta_{\tau}}+u_{\tau}))}{\sum_{t=1}^{\tau}\exp((z_{\beta_{t}}+u_{t})(z_{\beta_{\tau}}+u_{\tau}))}$
for all $(z,u)$, so when we multiply denominators and move everything to one side, we get that $\exp(1)$ is the root of a polynomial. Since $\exp(1)$ is transcendental, the polynomial is the zero polynomial, so each monomial has coefficient equal to zero. Thus, the idea is to choose $(z,u)$ such that the coefficient of a particular monomial being zero is precisely what we are trying to show. Moreover, we can do this for as many $(z,u)$’s as we want. First, we set $z=0$ and $u=(1,\ldots,T)$, targeting the lengths of $\alpha$ and $\beta$. Here, it makes sense that only the positional embedding is needed. Then, we set $z=(1,\ldots,\omega)$ and $u=\omega e_{s}$, using the positional embedding to target position $s$. These choices of $(z,u)$ turn out to be sufficient to show that $\alpha=\beta$.
Remark 6.2.
It is quite interesting that we needed the transcendence of $\exp(1)$ for the proof and it is natural to wonder what would happen if we replaced $\phi$ with the function $x\mapsto a^{x}/\sum_{\gamma=1}^{\omega}a^{x_{\gamma}}$ for some $a>1$ which is not transcendental. In this case, since $a^{b}=\exp(b\log(a))$, we would just need to scale our example $(z,u)$’s by $\sqrt{\log(a)}$ to get the same next-token prediction capacity result. So, the important aspect of self-attention that leads to its injectivity is that $\phi$ has the form $x\mapsto a^{x}/\sum_{\gamma=1}^{\omega}a^{x_{\gamma}}$, not that the base happens to be $\exp(1)$.
@@ 6.2 Rank results
$p(a)$ $\coloneqq\det_{I,J}\left(\psi\left(ab^{\top}\right)\right)=\sum_{S\in\binom{K}{s}}\left(\prod_{k\in S}c_{k}\right)\underset{\coloneqq p_{S}(a)}{\underbrace{\det_{I,[s]}\left(\left[a^{(k)}\right]_{k\in S}\right)}}\underset{\coloneqq q_{S}}{\underbrace{\det_{J,[s]}\left(\left[b^{(k)}\right]_{k\in S}\right)}}.$
@@
To see that $M$ is open, notice that it is the pre-image of $(-\rho,\rho)^{m\times n}$ under the continuous map $a\mapsto ab^{\top}$. Now, let $(c_{k})$ be the coefficients of the Taylor expansion at zero of $\psi$. Define $K=\{k\in\mathbb{N}\cup\{0\}\mid c_{k}\neq 0\}$. Define $s=\min\{m,n\}$ and let $k_{1},\ldots,k_{s}$ be the first $s$ integers in $K$. Define $K_{s}=\{k_{1},\ldots,k_{s}\}$. Let $I\subset[m]$ and $J\subset[n]$ such that $|I|=|J|=s$. Define $p:M\to\mathbb{R}:a\mapsto\det_{I,J}(\sum_{k\in K_{s}}c_{k}(ab^{\top})^{(k)})$ and $q:M\to R:a\mapsto\det_{I,J}(\psi(ab^{\top}))$. Note from the proof of Theorem 6.3 that the monomials of $p$ all have degree $k_{1}+\cdots+k_{s}$ whereas if $S\in\binom{K}{s}$ has any integers not in $K_{s}$, then its corresponding monomials will have degree $>k_{1}+\cdots+k_{s}$. Thus, $p$ is the truncation of the Taylor expansion at zero of $q$ to precisely the monomials of degree $k_{1}+\cdots+k_{s}$. Since, by Theorem 6.3, $p$ is not identically zero, it has at least one nonzero coefficient, and so the Taylor expansion at zero of $q$ has at least one nonzero coefficient. Thus, $q$ is not identically zero, completing the proof. ∎

### v3

@@ 5 Transformer Model
  • Let $Z\in\mathbb{R}^{d\times\zeta}$. This is the embedding matrix. Its $\gamma$th column corresponds to the embedding of token $\gamma$ in $\mathbb{R}^{d}$. We call $d$ the embedding dimension.
@@
  • Let $W\in\mathbb{R}^{d\times m},b\in\mathbb{R}^{m},V\in\mathbb{R}^{m\times\zeta}$. These are the hidden matrix, bias vector, and output matrix respectively. We call $m$ the number of neurons.
@@
$\begin{split}&h_{1,\theta}:\cup_{t=1}^{\infty}[\zeta]^{t}\to\cup_{t=1}^{\infty}\mathbb{R}^{d\times t}:\alpha\mapsto Z[:,\alpha]+U[:,1:|\alpha|],\\ &h_{2,\theta}:\cup_{t=1}^{\infty}\mathbb{R}^{d\times t}\to\mathbb{R}^{d}:X\mapsto\\ &\hskip 14.22636ptW_{0}^{\top}\text{vec}\left(\left[W_{3,r}^{\top}X\phi\left(X^{\top}W_{1,r}W_{2,r}^{\top}X[:,-1]\right)\right]_{r\in[m_{0}]}\right),\\ &h_{3,\theta}:\mathbb{R}^{d}\to\mathbb{R}^{\zeta}:x\mapsto V^{\top}\phi\left(W^{\top}x+b\right),\\ &\text{and }h_{\theta}\coloneqq\phi\circ h_{3,\theta}\circ h_{2,\theta}\circ h_{1,\theta}\end{split}$ (4)
where $-1$ denotes the last index and $\phi$ is the softmax function. We think of the value $h_{\theta}(\leavevmode\nobreak\ )$ as an additional $\zeta-1$ model parameters. The first three functions are the embedding, self-attention, and FNN sub-layers respectively of the one-layer transformer $h_{\theta}$. Note that we do not include the skip-connection $+X[:,-1]$ in the self-attention sub-layer. This was omitted for simplicity, but its inclusion does not affect Theorems 17 and 21, which follow by almost the same proof.
@@ 6 Transformer Next-token Prediction Capacity
In order to lower bound the next-token prediction capacity of the transformer model, we first reduce to the special case of $d=d_{1}=d_{0}=m_{0}=1$ and $W_{1,1}=W_{2,1}=W_{3,1}=W_{0}=1$. Thus, $h_{\theta}$ becomes the following composition of mappings:
@@
where $f(z,u,\alpha)=\langle x,\phi(xx_{|\alpha|})\rangle$ with $x=z[\alpha]+u[1:|\alpha|]$. To relate this to the previous section, $Z=z^{\top}$, $U=u^{\top}$, $W=w^{\top}$, and $h_{2,\theta}\circ h_{1,\theta}=f(z,u,\cdot)$. Assume $\psi:\mathbb{R}\to\mathbb{R}$ is real analytic at some $\eta\in\mathbb{R}$ and not a polynomial there. Assume $m\geq n$. We will show that this simplified model can interpolate arbitrary data sets of size $n$. Then we will use this to show that the full model can interpolate arbitrary data sets of size $n$.
@@
To show the former, let $\alpha^{1},\ldots,\alpha^{n}\in\cup_{t=1}^{\infty}[\zeta]^{t}$ and $Y^{\prime}\in(\mathrm{ri}\Delta^{\zeta-1})^{n}$. Then there exists $Y\in\mathbb{R}^{\zeta\times n}$ such that $\phi(Y)=Y^{\prime}$ with $\phi$ applied column-wise. We will construct an interpolating model. First, set $b=\eta\mathbbm{1}_{m}$. Second, define $T=\max_{i\in[n]}|\alpha^{i}|$ and sample $(z,u)\in\mathbb{R}^{\zeta}\times\mathbb{R}^{T}$ from any continuous probability distribution. Let $\hat{x}_{1},\ldots,\hat{x}_{n}$ be the corresponding outputs of the first mapping. Then the matrix of outputs of the composition of mappings is $\phi(V^{\top}\psi(w[\hat{x}_{i}]_{i\in[n]}^{\top}+b\mathbbm{1}_{n}^{\top}))$. In Theorem 17, we prove that the $\hat{x}_{i}$ are nonzero and distinct. In Theorem 20, we prove that this implies $\psi(w[\hat{x}_{i}]_{i\in[n]}^{\top}+b\mathbbm{1}_{n}^{\top})$ has full rank for generic $w$ in an open set around zero. So, as our third step in the construction, sample $w\in\mathbb{R}^{m}$ from any continuous probability distribution that has support in the open set. Then, $\psi(w[\hat{x}_{i}]_{i\in[n]}^{\top}+b\mathbbm{1}_{n}^{\top})$ has rank $n$ so, as our fourth and final step in the construction, solve for $V$ in $Y=V^{\top}\psi(w[\hat{x}_{i}]_{i\in[n]}^{\top}+b\mathbbm{1}_{n}^{\top})$ as a system of linear equations.
@@
Thus, the simplified model is able to interpolate arbitrary data sets as long as $m\geq n$. To see how to reduce the full model with $d,m_{0},d_{0}\in\mathbb{N}$ and $d_{r}\in\mathbb{N}\leavevmode\nobreak\ \forall r\in[m_{0}]$ to the simplified model, set $Z=\mathbbm{1}_{d}z^{\top}/\sqrt{d},U=\mathbbm{1}_{d}u^{\top}/\sqrt{d},W_{1,r}=W_{2,r}=\mathbbm{1}_{d}\mathbbm{1}_{d_{r}}^{\top}/\sqrt{dd_{r}}\leavevmode\nobreak\ \forall r\in[m_{0}],W_{3,r}=\mathbbm{1}_{d}\mathbbm{1}_{d_{0}}^{\top}/\sqrt{dd_{0}}\leavevmode\nobreak\ \forall r\in[m_{0}],W_{0}=\mathbbm{1}_{m_{0}d_{0}}\mathbbm{1}_{d}^{\top}/(m_{0}\sqrt{dd_{0}}),W=\mathbbm{1}_{d}w^{\top}/\sqrt{d}$. In other words, token $\gamma\in[\zeta]$ is embedded as $z_{\gamma}\mathbbm{1}_{d}/\sqrt{d}$ and the $t$th position vector is $u_{t}\mathbbm{1}_{d}/\sqrt{d}$. So, the inputs to the self-attention sub-layer are sequences of vectors coming from a one-dimensional subspace, and so it makes sense for the self-attention matrices to be constant matrices. Finally, the $k$th hidden neuron of the FNN sub-layer has weight vector $w_{k}\mathbbm{1}_{d}/\sqrt{d}$. By reducing to the simplified model in this way, and applying Theorems 17 and 20, we are able to lower bound the next-token prediction capacity of the one-layer multi-head decoder-only transformer model by $m$ in Theorem 21.
@@ 6.1 Injectivity of Self-attention
Theorem 17.
Let $\zeta,T\in\mathbb{N}$. Let $\phi$ be the softmax function. Define
$f:\leavevmode\nobreak$ $\mathbb{R}^{\zeta}\times\mathbb{R}^{T}\times\cup_{t=1}^{T}[\zeta]^{t}\to\mathbb{R}$
@@
First, let $A$ denote $\cup_{t=1}^{T}[\zeta]^{t}$. Then, note that, for all $z\in\mathbb{R}^{\zeta},u\in\mathbb{R}^{T},\tau\in[T],\alpha\in[\zeta]^{\tau}$,
@@
is real analytic. Note that the set of $(z,u)$ for which $f(z,u,\cdot)$ does not map injectively into \$\0\mathbb{R}\backslash\{0\}$ is precisely the zero set of $g$. Thus, the result follows if and only if $g$ is nontrivial if and only if (1) each $f(\cdot,\cdot,\alpha)$ is nontrivial and (2) each $f(\cdot,\cdot,\alpha)-f(\cdot,\cdot,\beta)$ is nontrivial. To prove (1), observe that $f(\mathbbm{1}_{\zeta},0,\alpha)=1\leavevmode\nobreak\ \forall\alpha\in A$. Now, we will go about proving (2). First, let $\alpha,\beta\in A$. We want to show that $\alpha\neq\beta$ implies $f(\cdot,\cdot,\alpha)-f(\cdot,\cdot,\beta)$ is nontrivial. We will prove the contrapositive. Suppose $f(\cdot,\cdot,\alpha)-f(\cdot,\cdot,\beta)\equiv 0$. First, we will show that $\alpha$ and $\beta$ have the same length, then we will show that each entry of $\alpha-\beta$ is zero.
@@
Let $\tau=|\alpha|=|\beta|$. Let $s\in[\tau]$. To show that $\alpha_{s}=\beta_{s}$, set $z=(1,\ldots,\zeta)$ and $u=\zeta e_{s}$, and define the polynomials
$p_{\gamma}=\sum_{t=1}^{\tau}(\gamma_{t}+\zeta\delta_{t,s})x^{(\gamma_{t}+\zeta\delta_{t,s})(\gamma_{\tau}+\zeta\delta_{\tau,s})}$
and $q_{\gamma}=\sum_{t=1}^{\tau}x^{(\gamma_{t}+\zeta\delta_{t,s})(\gamma_{\tau}+\zeta\delta_{\tau,s})}\leavevmode\nobreak\ \forall\gamma\in\{\alpha,\beta\}.$
Note that the monomial $x^{(\alpha_{s}+\zeta)(\alpha_{\tau}+\zeta\delta_{\tau,s})+(\beta_{s}+\zeta)(\beta_{\tau}+\zeta\delta_{\tau,s})}$ has coefficient $\alpha_{s}-\beta_{s}$ in $p_{\alpha}q_{\beta}-p_{\beta}q_{\alpha}$. Also, $f(z,u,\gamma)=p_{\gamma}(\exp(1))/q_{\gamma}(\exp(1))\leavevmode\nobreak\ \forall\gamma\in\{\alpha,\beta\}$. So, again, $p_{\alpha}q_{\beta}-p_{\beta}q_{\alpha}\equiv 0$. Thus, $\alpha_{s}=\beta_{s}$. Since this was for arbitrary $s\in[\tau]$, we get that $\alpha=\beta$, proving the theorem. ∎
@@
$\frac{\sum_{t=1}^{\tau}(z_{\alpha_{t}}+u_{t})\exp((z_{\alpha_{t}}+u_{t})(z_{\alpha_{\tau}}+u_{\tau}))}{\sum_{t=1}^{\tau}\exp((z_{\alpha_{t}}+u_{t})(z_{\alpha_{\tau}}+u_{\tau}))}$
$\hskip 14.22636pt=\frac{\sum_{t=1}^{\tau}(z_{\beta_{t}}+u_{t})\exp((z_{\beta_{t}}+u_{t})(z_{\beta_{\tau}}+u_{\tau}))}{\sum_{t=1}^{\tau}\exp((z_{\beta_{t}}+u_{t})(z_{\beta_{\tau}}+u_{\tau}))}$
for all $(z,u)$, so when we multiply denominators and move everything to one side, we get that $\exp(1)$ is the root of a polynomial. Since $\exp(1)$ is transcendental, the polynomial is the zero polynomial, so each monomial has coefficient equal to zero. Thus, the idea is to choose $(z,u)$ such that the coefficient of a particular monomial being zero is precisely what we are trying to show. Moreover, we can do this for as many $(z,u)$’s as we want. First, we set $z=0$ and $u=(1,\ldots,T)$, targeting the lengths of $\alpha$ and $\beta$. Here, it makes sense that only the positional embedding is needed. Then, we set $z=(1,\ldots,\zeta)$ and $u=\zeta e_{s}$, using the positional embedding to target position $s$. These choices of $(z,u)$ turn out to be sufficient to show that $\alpha=\beta$.
Remark 18.
It is quite interesting that we needed the transcendence of $\exp(1)$ for the proof and it is natural to wonder what would happen if we replaced $\phi$ with the function $x\mapsto a^{x}/\sum_{\gamma=1}^{\zeta}a^{x_{\gamma}}$ for some $a>1$ which is not transcendental. In this case, since $a^{b}=\exp(b\log(a))$, we would just need to scale our example $(z,u)$’s by $\sqrt{\log(a)}$ to get the same next-token prediction capacity result. So, the important aspect of self-attention that leads to its injectivity is that $\phi$ has the form $x\mapsto a^{x}/\sum_{\gamma=1}^{\zeta}a^{x_{\gamma}}$, not that the base happens to be $\exp(1)$.
@@ 6.2 Rank Results
$p(a)$ $\coloneqq\det_{I,J}\left(\psi\left(ab^{\top}\right)\right)$
$=\sum_{S\in\binom{K}{s}}\left(\prod_{k\in S}c_{k}\right)\underset{\coloneqq p_{S}(a)}{\underbrace{\det_{I,[s]}\left(\left[a^{(k)}\right]_{k\in S}\right)}}\underset{\coloneqq q_{S}}{\underbrace{\det_{J,[s]}\left(\left[b^{(k)}\right]_{k\in S}\right)}}.$
@@
To see that $M$ is open, notice that it is the pre-image of $(-\rho,\rho)^{m\times n}$ under the continuous map $a\mapsto ab^{\top}$. Now, let $(c_{k})$ be the coefficients of the Taylor expansion at zero of $\psi$. Define $K=\{k\in\mathbb{N}\cup\{0\}\mid c_{k}\neq 0\}$. Define $s=\min\{m,n\}$ and let $k_{1},\ldots,k_{s}$ be the first $s$ integers in $K$. Define $K_{s}=\{k_{1},\ldots,k_{s}\}$. Let $I\subset[m]$ and $J\subset[n]$ such that $|I|=|J|=s$. Define $p:M\to\mathbb{R}:a\mapsto\det_{I,J}(\sum_{k\in K_{s}}c_{k}(ab^{\top})^{(k)})$ and $q:M\to R:a\mapsto\det_{I,J}(\psi(ab^{\top}))$. Note from the proof of Theorem 19 that the monomials of $p$ all have degree $k_{1}+\cdots+k_{s}$ whereas if $S\in\binom{K}{s}$ has any integers not in $K_{s}$, then its corresponding monomials will have degree $>k_{1}+\cdots+k_{s}$. Thus, $p$ is the truncation of the Taylor expansion at zero of $q$ to precisely the monomials of degree $k_{1}+\cdots+k_{s}$. Since, by Theorem 19, $p$ is not identically zero, it has at least one nonzero coefficient, and so the Taylor expansion at zero of $q$ has at least one nonzero coefficient. Thus, $q$ is not identically zero, completing the proof. ∎